Pls help out: a number of two digits is increased by 9. When the digits are reversed, the sum of the digit is 5. Find the number
Answer by ankor@dixie-net.com(22740)

Two digits, a & b
then
10a+b = "the number"
:
Write an equation for each statement:
:
a number of two digits is increased by 9.
10a + b + 9
When the digits are reversed,
10a + b + 9 = 10b + a
10a - a + 9 = 10b - b
9a + 9 = 9b
divide by 9
a + 1 = b
a - b = -1
"the sum of the digit is 5"
a + b = 5
:
Use elimination
a - b = -1
a + b = 5
------------Adding eliminates b, find a
2a = 4
a = 2
Find b
2 + b = 5
b = 3
23 is the original number, add 9 you get 32, reversed digits, add up to 5